X-1 II. For f(x) find: x+3 Step 1: Analyze f(x) 4. Domain 5. X-intercepts and y-intercepts (use calculator to approximate valjust step 3 and 4
Answer #1

given f(x) = x-1 x+3 Quient Rale. Differentialing by f() = (2+3) d (x (x-1).d (x+3) dr da (x+3) ² f (x) = (x+3) thy (x-1) 2Sign chart of f(x) of -ve. the cocavacity C-a upward 3 (-as,-3) co uravity downward : (-3, co ) Step 4. graph 4 3 2 + 1 2 3 4